Determine whether each following statements about Fibonacci numbers is true or false




A. 2Fn >F n+1 for n≥3



B. 2F n+4 = f n+3 for n≥3


1
Expert's answer
2021-11-26T05:17:50-0500

The Fibonacci sequence, "F_0, F_1,F_2, \u2026 ," is defined by the initial condition "F_0=0," "F_1=1," and the recurrence relation "F_n = F_{n-1}+F_{n-2}" for "n=2, 3, 4, ..."


"0,1,1,2,3,5,8,13,21,34,55,89,144,..."

A.


"F_{n+1}=F_{n}+F_{n-1}<F_n+F_n=2F_n ,n\geq3"

The statement "2F_n>F_{n+1}" is true for "n\geq 3"


B.


"2F_{n+4}=2(F_{n+3}+F_{n+2})=F_{n+3}+F_{n+3}+2F_{n+2}"

Since "F_k>0" for "k\geq1," then


"2F_{n+4}=F_{n+3}+F_{n+3}+2F_{n+2}>F_{n+3}, n\geq3"

The statement "2F_{n+4}=F_{n+3}" is false for "n\geq 3."